Let ? and ? be random variables with finite means. Show that

min_{g(x)} E(Y-g(x))^{2} = E(Y-E(Y\mid X ))^{2}

Hint: ? − ? = ? − ? + ? − ?
Answer #1

X and Y be kus with Finite Meems ECY- g(x))2 = E{Y-/x) +E(4/X) – 4(x)}? - E (Y-E(Y/X)² +2 EY-EYRI + E{ECT (*) – g(x)}} = E(+Aence Eş y g(x)} 2 zx Mininom. IRF & holds FFF f(x) = E(47x) Hence min E&Y-9(x)} = E81-E (4) X)} - I proved] g(x)Note-if there is any understanding problem regarding this please feel free to ask via comment box..thank you